ok i need help with this problem:
1x+2y=2
-4x+3y=25
using the elimination process

Answer by checkley71(8403)

-4X+3Y=25 NOW MULTIPLY THE OTHER EQUATION BY 4 WE GET
4X+8Y=8 NOW ADD
--------------------------------------
11Y=33
Y=33/11
Y=3 SOLUTION
SUBSTITUTING 3 FOR Y WE GET
4X+8*3=8
4X+24=8
4X=-24+8
4X=-16
X=-16/4
X=-4 SOLUTION
PROOF
-4*-4+3*3=25
16+9=25
25=25